One number is 10 less than three times a second number. If their sum is 46, find the larger of the two numbers.
Answer by addingup(3677)

x = 3y-10
and:
x+y = 46 substitute for x:
3y-10+y = 46
4y = 56
y = 14 and
x = 3(14)-10 = 32
----------------------------
Check - the problem says:
x+y = 46
32+14 = 46 Correct.
:
John